# Vision Language Models Can Parse Floor Plan Maps

### v1

@@
Abstract
Vision language models (VLMs) can simultaneously reason about images and texts to tackle many tasks, from visual question answering to image captioning. This paper focuses on map parsing, a novel task that is unexplored within the VLM context and particularly useful to mobile robots. Map parsing requires understanding not only the labels but also the geometric configurations of a map, i.e., what areas are like and how they are connected. To evaluate the performance of VLMs on map parsing, we prompt VLMs with floorplan maps to generate task plans for complex indoor navigation. Our results demonstrate the remarkable capability of VLMs in map parsing, with a success rate of 0.96 in tasks requiring a sequence of nine navigation actions, e.g., approaching and going through doors. Other than intuitive observations, e.g., VLMs do better in smaller maps and simpler navigation tasks, there was a very interesting observation that its performance drops in large open areas. We provide practical suggestions to address such challenges as validated by our experimental results. Webpage: https://shorturl.at/OUkEY
I Introduction
A key to mobile robotics is a deep understanding of the geometric configuration of the world that mobile robots live in. As a result, many mobile robots need some forms of a map for localization, obstacle avoidance and navigation. To build such maps, the robots use a predefined data structure, e.g., an occupancy grid [1] or visual features [2], and then perform simultaneous localization and mapping (SLAM). Human beings have a long history of building and using maps. These days one can easily read floorplan maps of airports and shopping centers, and figure out a plan for navigation. By comparison, robots can hardly reach comparable competency in map reading and task planning. It is a non-trivial task for the robots due to the many labels in the map, their ambiguous associations to different areas, and complex geometric configurations. As a result, there is no existing method for addressing the map parsing problem, i.e., computing a navigation plan given a map image and a goal text, which motivated this research.
@@ I Introduction
Vision language models (VLMs) are foundation models that learn from and reason about both images and texts, supporting a variety of downstream tasks from visual question answering to image captioning. VLMs have demonstrated impressive successes in a variety of applications [3, 4, 5] including those in robotics [6, 7]. We, for the first time, apply VLMs to the novel task of map parsing and evaluate its performance in navigation, a foundational problem in mobile robotics. Our approach is simple and intuitive. A floorplan map and a problem description, including the start and goal positions, are provided to a VLM, and the task is to compute a plan (i.e., an action sequence) to achieve the goal. Despite the straightforward idea, the results are surprisingly good. Navigation plans generated by VLMs which can require a sequence of nine actions are generated correctly up to 90% of the time on some floorplans.
@@
II Related Work
II-A Existing Map Representations and Navigation
Existing works demonstrating autonomous navigation on mobile robots usually require generating an occupancy grid [8, 9, 10, 11], or leveraging vision-based methods for simultaneous localization and mapping (Visual SLAM) [2, 12, 13, 14, 15]. While such map representations have proven to be effective for autonomous navigation tasks, generating an accurate map is oftentimes labor-intensive. For instance, in vision-based settings, the robot has limited knowledge of the global environment, either leading to lengthy exploration, or navigational commands from a human. In this work, we greatly reduce the effort of generating accurate maps while still leveraging detailed information of the environment through the use of existing, and potentially in situ, floor plans.
II-B VLM Prompting Strategies
The output of a large pre-trained model largely relies on the way it is prompted. Strategies like chain-of-thought [16] and in-context learning [17] are leveraged on LLMs to improve performance. Similar to language prompts, image prompting strategies can improve VLM outputs. Set-of Mark prompting, which segments and labels objects in an image has shown to improve VLM responses [18]. In our work, we design a new visual prompting strategy designed for obtaining a spatial understanding of floor plan images.
II-C Large Models in Robotics
To improve the common-sense reasoning capabilities of robots, large pre-trained LLMs and VLMs have been integrated in robots for various tasks like housekeeping [19], object rearangement [20], navigation [21, 22, 23, 24, 25], and quadruped locomotion [26, 27] to name a few. Alignment of the large model with the environment and robot’s skills is critical to perform tasks in the real world [28]. Various approaches have demonstrated planning capabilities [29, 30] and uncertainty estimation [31] of large models. Various visual prompting strategies designed for different robotic manipulation tasks have also been developed [6, 7]. In line with recent works that leverage large models to incorporate common-sense knowledge on robots, we generate feasible navigation plans directly from an image of a floor plan.
III Methodology
III-A Visual Prompting Strategy
Our study uses floor plan images that include detailed architectural layouts for various building types. To generate accurate plans from a VLM, it’s important to design a visual prompt which can facilitate learning the structure of the floor plan. Unfortunately, raw floor plan images tend to contain various markings (i.e. windows, furniture symbols, non-uniform wall thickness) which can potentially confuse the VLM in understanding the general layout of the floor plan. Thus, we remove such markings to produce a cleaner map which can be better leveraged by a VLM.
We find that removing extraneous details from the floor plan is insufficient for the VLM to understand the map layout. In particular, we find the VLM has limited spatial awareness for sparsely labelled rooms with lots of open space, and near key decision points like doors and intersections. To alleviate this, we add duplicate room labels in open spaces and near doors and intersections. This provides the VLM with further guidance in understanding the general structure of the floor plan. We later demonstrate the importance of such additional labels in Section IV. In this work, we manually remove extraneous markings and add additional room labels for visual prompt construction. We leave automated floor plan editing as future work.
Map enhancement with dense labeling: A key finding in this research is the importance of dense labeling in map enhancement. We apply a methodology in which rooms are labeled strategically at decision points (e.g., near doors or intersections). We later show that such label placement brings a considerable performance boost, which is particularly significant in complex environments where navigation paths can involve multiple rooms and transitions. This process can be automated (though not in this paper) to make it scalable for broader applications without manual intervention.
III-B VLM-Based Plan Generation
In our VLM-based framework, as shown in Fig. 2, VLMs formulate navigation plans based on a floor plan image and a text prompt. This method leverages an instruction-based text prompting strategy, which involves providing the VLM with explicit and detailed guidelines for the navigation task, along with the floor plan image. We now elaborate on the prompting strategy and the resulting output format.
The text prompt given to the VLM is shown in Fig. 3. The prompt explicitly defines the starting point and the destination. This allows the VLM to understand the required navigation path and objectives clearly. It provides detailed instructions on interpreting the floor plan and managing door interactions. These guidelines include specific protocols for door operations and decision-making processes.
@@ III-B VLM-Based Plan Generation
Figure 5: Example of a doubled map.
(a) Map Size
(b) Task Difficulty
(c) Label Density
Figure 6: Comparison of GPT-4o results for map size, task difficulty and label density. We used original maps in the “label density” experiment. To accommodate hard tasks, we used doubled maps in the “task difficulty” experiment.
(a) Map Size
(b) Task Difficulty
(c) Label Density
Figure 7: Comparison of Claude Sonnet 3.5 results for map size, task difficulty and label density. We used original maps in the “label density” experiment. To accommodate hard tasks, we used doubled maps in the “task difficulty” experiment.
IV Experiments
IV-A Experimental Setup
Our study uses floor plan images from a publicly available dataset CVC-FP [32], which includes detailed architectural floor plans for various building types. Three floor plans were randomly selected from those that contain 9-11 rooms and clear labels. Those maps are referred to as “original maps” and are shown in Fig. 4. Our experiments use two state-of-the-art VLMs: GPT-4o [33] and Claude-3.5 Sonnet [34].
@@ IV-A Experimental Setup
IV-A1 Map Size
To examine the impact of map size on navigation performance, we developed two types of maps: Original Maps that are enhanced versions of the maps selected from the CVC-FP dataset, and Doubled Maps that were created by connecting two copies of an original map through an additional door. Fig. 5 presents an example of a doubled map. A door denoted as D10 is added to establish connectivity, thereby forming the final doubled map.
IV-A2 Task Difficulty
@@
IV-A3 Label Density
We evaluated the impact of label density on model performance by implementing two labeling schemes. Sparse-Labeled maps are those where each room was labeled with a single label, usually placed at the center. This minimalistic approach offered fewer cues for the model to base its navigation decisions on. Dense-Labeled maps include multiple labels for each room, where the placement is described in Section III-A.
IV-B Experimental Results
IV-B1 Hypothesis 1 (Map Size)
@@
The results indicate that accuracy decreases as map size increases in the overall domain analysis, with the VLMs performing better in smaller maps. The difference is significant, as a T-test on trials with GPT-4o revealed a drop in accuracy ($t=6.13$, $p<0.0001$). This supports our hypothesis that accuracy decreases as map size increases. The results are reported in Fig. 6 and 7. Both GPT and Claude exhibit a similar pattern of performance decline with larger maps.
IV-B2 Hypothesis 2 (Task Difficulty)
@@
The accuracy of the GPT-4o models generally decreased with increased task difficulty, as more complex tasks led to lower accuracy in two out of the three maps. For example, a T-test on doubled_map_2 indicated a drop in accuracy with $t=2.88$ and $p=0.0047$. The results from both VLMs generally support our hypothesis that more complex tasks lead to lower accuracy, as navigating through multiple rooms introduces additional challenges. One example task is detailed in Table I.
TABLE I: An example navigation plan generated by GPT-4o in Map 1 shown on the left of Fig. 4. The initial location is “Terrrasse Couverte” and the goal location is “Chambre 1”. This navigation task requires a sequence of nine actions. GPT-4o achieved 0.96 success rate in this map on similar hard navigation tasks, which demonstrates great promises for VLM-based map parsing research.
@@ 9 GoThrough(D4)
Figure 8: The robot localizes directly on the floor plan image and performs a navigation task. The occupancy grid map that the robot used for localization and navigation was also derived from the floor plan map captured by the robot. We manually removed the labels that would otherwise be interpreted by the robot as obstacles. From right to left, the robot was performing 1) Going through Door N09, 2) Approaching Door N00, 3) Opening Door N00, and 4) Approaching Door T01.
IV-B3 Hypothesis 3 (Label Density)
@@
V Hardware Demonstration
Our VLM-based planning and navigation system is demonstrated on a DEEPRobotics Lite3 quadruped robot. An image of the floor plan of a building on a college campus is captured from the robot’s camera. This image is then edited as described in Section III to make it suitable for the VLM query. Another version of the raw image with space whited out is directly used for robot localization, as seen in Fig. 8. The VLM is then queried with the edited photo, and a navigation plan consisting of a sequence of navigation and door opening actions is generated. The robot executes this navigation plan, to move from the robotics lab (room N09), to a classroom (room T01). The robot localizes itself directly on a grayscale version of the floor plan, without the need for generating an accurate occupancy grid via SLAM. The robot successfully avoids the obstacles not present on the floor plan, asks for doors to be opened as needed, and achieves the desired navigation goal.
VI Conclusion and Future Work

### v2

# footnotetext:

@@
Abstract: Vision language models (VLMs) can simultaneously reason about images and texts to tackle many tasks, from visual question answering to image captioning. This paper focuses on map parsing, a novel task that is unexplored within the VLM context and particularly useful to mobile robots. Map parsing requires understanding not only the labels but also the geometric configurations of a map, i.e., what areas are like and how they are connected. To evaluate the performance of VLMs on map parsing, we prompt VLMs with floor plan maps to generate task plans for complex indoor navigation. Our results demonstrate the remarkable capability of VLMs in map parsing, with a success rate of 0.96 in tasks requiring a sequence of nine navigation actions, e.g., approaching and going through doors. Other than intuitive observations, e.g., VLMs do better in smaller maps and simpler navigation tasks, there was a very interesting observation that its performance drops in large open areas. We provide practical suggestions to address such challenges as validated by our experimental results. Webpage: https://sites.google.com/view/vlm-floorplan/.
Keywords: task planning; mobile robots; vision language models; robot navigation
1 Introduction
A key to mobile robotics is a deep understanding of the geometric configuration of the world that mobile robots live in. As a result, many mobile robots need some forms of a map for localization, obstacle avoidance and navigation. To build such maps, the robots use a predefined data structure, e.g., an occupancy grid elfes1989using or visual features taketomi2017visual , and then perform simultaneous localization and mapping (SLAM). Human beings have a long history of building and using maps. These days one can easily read floor plan maps of airports and shopping centers, and figure out a plan for navigation. By comparison, robots can hardly reach comparable competency in map reading and task planning. It is a non-trivial task for the robots due to the many labels in the map, their ambiguous associations to different areas, and complex geometric configurations. As a result, there is no existing method for addressing the map parsing problem, i.e., computing a navigation plan given a map image and a goal text, which motivated this research.
@@ 1 Introduction
Vision language models (VLMs) are foundation models that learn from and reason about both images and texts, supporting a variety of downstream tasks from visual question answering to image captioning. VLMs have demonstrated impressive successes in a variety of applications antol2015vqa , chen2020uniter , brooks2023instructpix2pix . Among them, robotics is an important application domain nasiriany2024pivot , liu2024moka , hu2023toward . We, for the first time, apply VLMs to the novel task of map parsing and evaluate its performance in navigation, a foundational problem in mobile robotics. Our approach is simple and intuitive. A floor plan map and a problem description, including the start and goal positions, are provided to a VLM, and the task is to compute a plan (i.e., an action sequence) to achieve the goal. Figure 1 shows a quadruped mobile robot executing a navigation plan that was generated by a VLM using a floor plan map. Despite the straightforward idea, the results are surprisingly good. Navigation plans generated by VLMs which can require a sequence of nine actions are generated correctly up to 90% of the time on some floor plans.
@@
2 Related work
2.1 Existing map representations and navigation
Existing works demonstrating autonomous navigation on mobile robots can be categorized into two groups. One relies on generating an occupancy grid map zhang2015mobile , hayamizu2021guiding , fu2022coupling , defazio2023seeing , where each pixel in the map identifies an area being occupied, obstacle-free or unknown. Another popular way of realizing mobile navigation leverages vision-based methods for simultaneous localization and mapping (Visual SLAM) taketomi2017visual , zhang2021frl , macario2022comprehensive , sorokin2022learning , hwang2023system . Those methods rely on visual features, and then apply simultaneously estimate the locations of both visual landmarks and the robot itself thrun2002probabilistic . While such map representations have proven to be effective for autonomous navigation tasks, generating an accurate map is oftentimes labor-intensive. For instance, in vision-based settings, the robot has limited knowledge of the global environment, either leading to lengthy exploration, or navigational commands from a human. In this work, we greatly reduce the effort of generating accurate maps while still leveraging detailed information of the environment through the use of existing, and potentially in situ, floor plans.
2.2 VLM prompting strategies
The output of a large pre-trained model largely relies on the way it is prompted. Strategies like chain-of-thought wei2022chain , in-context learning dong2022survey and others yao2023tree , besta2024graph are leveraged on LLMs to improve performance. Similar to language prompts, image prompting strategies can improve VLM outputs. Set-of Mark prompting, which segments and labels objects in an image has shown to improve VLM responses yang2023set . In our work, we design a new visual prompting strategy designed for obtaining a spatial understanding of floor plan images.
2.3 Large models in robotics
To improve the common-sense reasoning capabilities of robots, large pre-trained LLMs and VLMs have been integrated into robotic systems for various tasks. Housekeeping kant2022housekeep , li2023behavior and object rearrangement ding2023task , huang2019large , danielczuk2021object are two of the service task examples. Navigation is another robotics domain where LLMs and VLMs have demonstrated exceptional performance yokoyama2024vlfm , rajvanshi2024saynav , song2024socially , sathyamoorthy2024convoi , chen2024commonsense . Furthermore, while quadruped locomotion problems traditionally were solved using optimization methods, the effectiveness of LLMs and VLMs has been evident in recent years yu2023language , tang2023saytap , gu2025humanoid . Alignment of the large model with the environment and robot’s skills is critical to perform tasks in the real world ahn2022can . Various approaches have demonstrated planning capabilities huang2022inner , liu2023llm+ and uncertainty estimation ren2023robots of large models. Various visual prompting strategies designed for different robotic manipulation tasks have also been developed nasiriany2024pivot , liu2024moka . In line with recent works that leverage large models to incorporate common-sense knowledge on robots, we generate feasible navigation plans directly from an image of a floor plan.
3 Methodology
3.1 Visual prompting strategy
Our study uses floor plan images that include detailed architectural layouts for various building types. To generate accurate plans from a VLM, it’s important to design a visual prompt which can facilitate learning the structure of the floor plan. However, raw floor plan images tend to contain various markings (i.e. windows, furniture symbols, non-uniform wall thickness) which can potentially confuse the VLM in understanding the general layout of the floor plan. Thus, we remove such markings to produce a cleaner map which can be better leveraged by a VLM.
We find that removing extraneous details from the floor plan is insufficient for the VLM to understand the map layout. In particular, we find the VLM has limited spatial awareness for sparsely labelled rooms with lots of open space, and near key decision points like doors and intersections. To alleviate this, we add duplicate room labels in open spaces and near doors and intersections. This provides the VLM with further guidance in understanding the general structure of the floor plan. We later demonstrate the importance of such additional labels in Section 4
In this study, all floor plans are preprocessed through a simple yet systematic enhancement procedure, Next, we present the step-by-step instructions that we followed for map enhancement.
  1. (1)
    Visual simplification: Raw floor plans frequently include architectural annotations outside the map itself, such as dimensions and scale bars. We remove those visuals and keep only the map part of the floor plans. In addition, there are non structural visuals, such as window icons, and furniture, where people use very different icons or symbols. Those were removed for standardizing evaluations.
  2. (2)
    Structural standardization: The line segments for walls and doors can be different in floor plan maps. We adjust the wall thickness to a uniform value. Doors are colored yellow, making their locations visually salient and uniformly represented across maps. Windows are replaced with walls to avoid confusing the robots on whether windows are navigable.
  3. (3)
    Dense labeling: Some rooms are very large, of irregular shape (e.g., long corridors), or both. We added additional labels of rooms at areas near doors or intersections, or within large open areas. This dense labeling provides additional contextual cues, leading to significant performance improvements, especially in complex environments involving multiple rooms and transitions.
  4. (4)
    Consistency check: The last step is to ensure that all rooms are enclosed, no overlapping or misleading elements remain, and every object (rooms and doors in particular) has a unique ID. The finalized map is then exported as a high-resolution PNG image as the image prompt of the VLMs.
In this work, the map enhancement steps were manually performed, though the process is straightforward and can be reproduced by robotics practitioners using our instructions. We discuss possible ways of automating the map enhancement process near the end of this article as future work.
3.2 VLM-based plan generation
In our VLM-based framework, as shown in Figure 2, VLMs formulate navigation plans based on a floor plan image and a text prompt. This method leverages an instruction-based text prompting strategy, which involves providing the VLM with explicit and detailed guidelines for the navigation task, along with the floor plan image. We now elaborate on the prompting strategy and the resulting output format.
@@ 3.2 VLM-based plan generation
The text prompt given to the VLM is shown in Figure 3. The prompt explicitly defines the starting point and the destination. This allows the VLM to understand the required navigation path and objectives clearly. It provides detailed instructions on interpreting the floor plan and managing door interactions. These guidelines include specific protocols for door operations and decision-making processes.
@@
4 Experiments
4.1 Experimental setup
Our study uses floor plan images from a publicly available dataset CVC-FP Heras15a , which includes detailed architectural floor plans for various building types. Three floor plans were randomly selected from those that contain 9–11 rooms and clear labels. Those raw maps are shown in Figure 4. As illustrated, these images often contain irrelevant visual elements such as scale bars, dimension lines, and furniture/window symbols, which are not pertinent to the general layout. To facilitate more effective visual prompting for VLMs, we preprocess these raw maps using the procedure described in Section 3.1 The resulting maps are referred to as “original maps” and are shown in Figure 5. Our experiments use two state-of-the-art VLMs: GPT-4o GPT-4o and Claude-3.5 Sonnet Claude35 .
Figure 4: Three raw floor plan maps.
@@ 4.1 Experimental setup
4.1.1 Map size
To examine the impact of map size on navigation performance, we developed two types of maps: Original Maps that are enhanced versions of the maps selected from the CVC-FP dataset, and Doubled Maps that were created by connecting two copies of an original map through an additional door. Figure 6 presents an example of a doubled map. A door denoted as D10 is added to establish connectivity, thereby forming the final doubled map.
Figure 6: Example of a doubled map.
4.1.2 Task difficulty
@@
4.1.3 Label density
We evaluated the impact of label density on model performance by implementing two labeling schemes. Sparse-Labeled maps are those where each room was labeled with a single label, usually placed at the center. This minimalistic approach offered fewer cues for the model to base its navigation decisions on. Dense-Labeled maps include multiple labels for each room, where the placement is described in Section 3.1
4.2 Experimental results
4.2.1 Hypothesis 1 (map size)
@@
The results indicate that accuracy decreases as map size increases in the overall domain analysis, with the VLMs performing better in smaller maps. The difference is significant, as a T-test on trials with GPT-4o revealed a drop in accuracy ($t=6.13$, $p<0.0001$). This supports our hypothesis that accuracy decreases as map size increases. The results are reported in Figures 7 and 8. Both GPT and Claude exhibit a similar pattern of performance decline with larger maps.
Figure 7: Comparison of GPT-4o results across three experimental conditions: (a) Map size: performance on original maps vs. doubled maps over hard tasks; (b) Task difficulty: performance on doubled maps over easy vs. hard tasks; (c) Label density: performance on original maps annotated with sparse vs. dense labels over hard tasks.
Figure 8: Comparison of Claude Sonnet 3.5 results across three experimental conditions: (a) Map Size: performance on original maps vs. doubled maps over hard tasks; (b) Task Difficulty: performance on doubled maps over easy vs. hard tasks; (c) Label Density: performance on original maps annotated with sparse vs. dense labels over hard tasks.
When the map size is enlarged, it was observed that the VLM agents frequently had difficulties in figuring out door ownership. For instance, some of the generated plan included infeasible actions such as going through door D from room R, where the door (D) does not belong to room R. This is a strong indication that in large floor plans, it is challenging for the VLMs to reason about door ownership. By comparison, such mistakes were rarely seen in small floor plans.
We looked into the navigation plans generated by GPT-4o for hard tasks in the three original maps reported in Figure 7, where dense labels were provided. Each map contained five tasks, and each task was run ten times. For instance, Domain 3 included the five navigation tasks of (‘GARAGE’, ‘BUREAU’), (‘LINGERIE’, ‘GARAGE’), (‘GARAGE’, ‘CHAMBRE ENFANT 1’), (‘CHAMBRE PARENTS’, ‘GARAGE’), and (‘GARAGE’, ‘WC’). Among the total 150 trials, GPT-4o was successful in 135 trials by generating navigation plans that correctly connected the initial and goal rooms. Among those 135 successful trials, 132 were optimal. In the 10 trials of the task (‘GARAGE’, ‘WC’), the VLM failed in one trial, and reported suboptimal plan D4–D3–D2–D1 three times, while the optimal path D4–D11–D1 was reported six times. For all other successful trials, the reported plans were optimal.
4.2.2 Hypothesis 2 (task difficulty)
@@
The accuracy of the GPT-4o models generally decreased with increased task difficulty, as more complex tasks led to lower accuracy in two out of the three maps. For example, a T-test on doubled_map_2 indicated a drop in accuracy with $t=2.88$ and $p=0.0047$. The results from both VLMs generally support our hypothesis that more complex tasks lead to lower accuracy, as navigating through multiple rooms introduces additional challenges. One example task is detailed in Table 1.
Table 1: An example navigation plan generated by GPT-4o in Map 1 shown on the left of Figure 5. The initial location is “Terrrasse Couverte” and the goal location is “Chambre 1”. This navigation task requires a sequence of nine actions. GPT-4o achieved 0.96 success rate in this map on similar hard navigation tasks, which demonstrates great promises for VLM-based map parsing research.
@@ 9 GoThrough(D4)
Difficult tasks required the VLMs to generate navigation plans to go through three doors, which required strong reasoning capabilities about room connectivity and graph topology. One typical failure case was that VLMs insisted on directly entering a room between the start and goal rooms, when there is no direct access to the room in the middle. Consider the floor plan in Figure 9a. When the robot started in room “N09” and aimed to go to room “Stair 3”, VLMs frequently suggested the robot to directly enter room N01. While N01 is immediately adjacent to both N09 and Stair 3, the robot cannot directly enter N01 from N09, because the only door of N01 connects to Hall2 instead of N09. Such observations indicate that VLMs had difficulty in reasoning about topological graphs of room connectivity through doors.
Figure 9: Three additional maps from the college campus. (a) Engineering Building; (b) Administration Building; (c) Library.
4.2.3 Hypothesis 3 (label density)
@@
Without dense labeling, VLMs frequently made mistakes in reasoning about door ownership and room connectivity through doors. For instance, door D16 in Figure 9b connects two rooms “NORTH READING ROOM” and “HALLWAY”, where the former is very large. We found that the VLMs sometimes wrongly believed that D16 connects “MEN RESTROOM” and “HALLWAY” because those two room labels are closer to the door label. To facilitate VLMs to reason about such door ownerships, we added two additional labels for the reading room, and one of them was placed next to door D16, where dense labeling was found useful for map parsing and navigation task planning.
@@
4.2.4 Raw vs. processed floor plan maps
We evaluated VLM performance on raw floor plan maps for the hard tasks. As shown in Figure 4, these raw images contain various irrelevant elements that are not essential to the spatial layout, such as furniture symbols, dimension lines, and scale bars. The results, presented in Table 2, show that accuracy is substantially lower when using raw maps compared to the processed sparse-label maps. This supports our claim in Section 3.1 that raw maps contain distracting visual elements which may confuse the VLM in understanding the overall layout.
Table 2: Average accuracy results comparing processed and raw floor plan maps across three domains for hard tasks.
Map Type GPT-4o Claude Sonnet 3.5
Domain 1 Domain 2 Domain 3 Domain 1 Domain 2 Domain 3
Processed Maps 0.24 0.84 0.22 0.68 0.66 0.34
Raw Maps 0.03 0.21 0.05 0.08 0.18 0.22
4.2.5 Diverse floor plans
To evaluate the generalizability of our conclusions across different environments, we extended our experiments to three additional floor plans selected from the home institution of this article’s authoring team, including the Engineering Building, the Library, and the Administration Building. The Engineering Building is also the environment where we conducted the real-world demonstration, described in Section 5 on Hardware Demonstration. These three floor plans are illustrated in Figure 9. The results, presented in Table 3, show that the VLMs performed well on the processed maps.
Table 3: Average accuracy of VLM performance on three college campus floor plans.
Task Difficulty GPT-4o Claude Sonnet 3.5
Engineering Administration Library Engineering Administration Library
Easy Task 0.92 0.90 0.92 0.78 0.70 0.94
Hard Task 0.80 0.94 0.90 0.54 0.62 0.90
5 Hardware demonstration
Our VLM-based planning and navigation system is demonstrated on a DEEP Robotics Lite3 quadruped robot. An image of the floor plan of a building on a college campus is captured from the robot’s camera. This image is then edited as described in Section 3 to make it suitable for the VLM query. Another version of the raw image with space whited out is directly used for robot localization, as seen in Figure 10. The VLM is then queried with the edited photo, and a navigation plan consisting of a sequence of navigation and door opening actions is generated. The robot executes this navigation plan, to move from the robotics lab (room N09), to a classroom (room T01). The robot localizes itself directly on a grayscale version of the floor plan, without the need for generating an accurate occupancy grid via SLAM.
Figure 10: The robot localizes directly on the floor plan image and performs a navigation task. The occupancy grid map that the robot used for localization and navigation was also derived from the floor plan map captured by the robot. We manually removed the labels that would otherwise be interpreted by the robot as obstacles. From right to left, the robot was performing (1) Going through Door N09, (2) Approaching Door N00, (3) Opening Door N00, and (4) Approaching Door T01.
The robot’s software for localization, obstacle avoidance and navigation, including actions of approaching and going through doors, was constructed using an existing software codebase for building-wide robot intelligence khandelwal2017bwibots . The codebase supported labeling door areas on both sides of a door. Whenever the robot gets inside either of the door areas, it is believed that the robot successfully completes the action of approaching the door. The action of going through a door is realized by moving from one of the door areas to the other side of the same door. Our robot cannot open doors by itself and door opening actions are realized by repeatedly saying “Please help me open the door” to seek human help until the two areas of the door become accessible to each other. We used the Lite3 robot’s locomotion skills that rely on its default Model Predictive Controller (MPC). The robot used standard methods from the ROS community for navigation, including the dynamic window approach (DWA) for obstacle avoidance, A* for global path planning, and Adaptive Monte Carlo Localization (AMCL) for localization using grayscale occupancy-grid map images. The robot successfully avoided the obstacles, asked for doors to be opened as needed, and achieved the desired navigation goal, as illustrated in Figure 10.
6 Conclusion and future work